A mineral contains 14.7% calcium and tungsten the remainder is oxygen find its emperical formula


1
Expert's answer
2021-07-29T23:02:01-0400

"Oxygen = 67.5-14.7= 52.8"


Moles

"calcium =\frac{14.7}{40.08}=0.367"


"tungsten =\frac{67}{183.84}=0.367"


"Oxygen =\frac{52.8}{32}=1.65"


Mole ratio


"Calcium =\frac {0.367}{0.367}=1"


"Tungsten =\frac {0.367}{0.367}=1"


"Oxygen =\frac {1.65}{0.367}=4.49= 4"


Empirical formula CaO4W